# Probing soft signals of gravitational-wave memory with space-based interferometers

### 30 Mar 2026 v1

@@ Abstract
Gravitational-wave displacement memory is a remarkable and ubiquitous phenomenon predicted by general relativity, which has not yet been detected. Unlike the oscillatory components of gravitational waveforms, displacement memory is associated with soft gravitons, making it the only observable signal of its parent event at sufficiently low frequencies. Similarly, soft waveforms may arise from velocity and integrated-displacement memory. The simple and universal spectral shapes of soft waveforms also provide effective templates for matched filtering and parameter estimation. In this paper, we systematically investigate the detection prospects for such soft memory signals with future space-based laser interferometers. As realistic examples, we examine the infrared spectral features of gravitational waves from moderately relativistic compact binary scattering and quasi-circular, non-precessing black hole mergers. In both cases, the low-frequency spectrum can be described by a soft waveform of displacement memory with a real correction factor. The results of simulated Bayesian parameter estimation demonstrate that independent measurement of a soft displacement-memory signal with a single LISA-like detector is achievable at signal-to-noise ratios $\gtrsim 10$. The measurement precision can be significantly improved by joint observations with a LISA-Taiji network. A single BBO detector would be capable of separately measuring the null memory from stellar-mass compact binary mergers. We also evaluate the detectability of an idealized stochastic background of soft displacement-memory signals. Our results indicate that gravitational-wave bursts with memory can be promising targets for space-based interferometers.
@@ I Introduction
Besides the oscillatory components, gravitational radiation in general relativity generally contains a non-oscillatory component that results in a displacement memory [10, 18, 85, 31]. It may have an “ordinary” or “null” origin [8], and is associated with the supertranslation charge balance laws of asymptotically flat spacetimes [91, 33, 78, 67] as well as with Weinberg’s soft graviton theorem [78]. Displacement memory can therefore serve as an important test of general relativity. In addition to their theoretical interest, gravitational wave (GW) bursts with displacement memory are expected to be produced by a variety of astrophysical sources, including compact binaries [30, 29, 32, 84, 94, 26, 87, 63, 100, 55, 47, 96, 24, 44, 64, 38, 48, 28, 22, 5], gamma-ray bursts [75, Yu_2020, 76], and core-collapse supernovae [68, 72]. Searches for displacement memory in the GWs from compact binary mergers have been considered for ground-based detectors [62, 66, 94, 27, 53, 52, 87, 42, 17, 41], space-based detectors [56, 40, 39, 54, 81, 41, 19, 101], and pulsar timing arrays [88, 89, 56, 2, 79, 25, 80, 1]. Potentially observable imprints of displacement memory may also appear in the cosmic microwave background [102].
@@
Although displacement memory ubiquitously occurs in gravitational radiation, it has so far remained undetected. For example, the null memory of a compact binary merger produces only a small modification of the waveform and is not easily disentangled from the oscillatory “parent” signal in the time domain. Resolving such modifications through sensitive detectors and refined waveform modeling would enable the detection of memory and the improvement of parameter estimation [42, 39, 54, 81]. Accurately modeling waveforms that include memory effects becomes increasingly challenging for more complex sources. For a compact binary scattering event, even with a sufficiently accurate waveform model, parameter estimation is still challenging due to degeneracies [fontbuté2024numericalrelativitysurrogatemodelhyperbolic]. Waveform modeling is further complicated by possible environmental influences [15, 77]. However, because of its soft nature, displacement memory produces a signal that persists at low frequencies, with a spectrum of universal shape. Although much information about the sources is lost in such “orphan” signals [88, 66, 40], they remain valuable targets for probing the memory components—either by searching for them or by measuring them independently—thanks to the ubiquity of the sources and the simplicity of the templates. It is also possible that some yet-unknown processes could produce observable GW bursts with displacement memory, making soft memory signals a potential probe of these processes.
@@
In addition to the soft waveform associated with displacement memory (which corresponds to a step-like strain in the time domain), other types of soft waveforms may also be observed, including those associated with velocity memory [43, 9] (corresponding to a step-like first derivative of strain in the time domain) and integrated-displacement memory (discussed in Sec. II.3). Although these waveforms are not related to asymptotic symmetries or the soft theorem, they can arise over sufficiently long time scales and are likewise associated with soft gravitons.
@@
In this paper, we study the prospects for detecting and measuring such soft memory signals with future space-based interferometers, including Laser Interferometer Space Antenna (LISA) [3, 4], Taiji [50], TianQin [50], and Big Bang Observer (BBO) Stage I [20]. We investigate the detector responses and the parameter estimation precision for both a single LISA-like detector and the proposed LISA-Taiji network [74]. In optimal configurations, the signal-to-noise ratio in LISA, Taiji and TianQin can reach 10 for a displacement memory of amplitude $\sim 10^{-20}$, or a velocity memory of amplitude $\sim 10^{-22}\,\text{Hz}$. To examine the precision of the measurement, we perform numerical simulations for Bayesian parameter estimation. We find that with a single detector, the memory parameters can be effectively probed for signal-to-noise ratios $\gtrsim 10$, although the measurement precision is limited by the signal’s degeneracy at low frequencies. This limitation can be mitigated by using a detector network or by incorporating prior information on the source direction. We demonstrate that the joint observations by LISA and Taiji can achieve significantly more precise measurements. Soft memory signals in BBO (occurring at $f\lesssim 1\,\text{Hz}$) appear to suffer from less degeneracy and can achieve much higher sensitivity, with the signal-to-noise ratio reaching 10 for a displacement memory of amplitude $\sim 10^{-24}$, or a velocity memory of amplitude $\sim 10^{-23}\,\text{Hz}$. This makes it possible to independently measure the null displacement memory from stellar-mass compact binary mergers.
@@
The soft waveform is an idealized zero-frequency limit. We examine the correction to the soft waveform in two realistic examples: hyperbolic encounter of compact binary and quasi-circular BBH merger, and assess its impact on displacement memory measurements using the uncorrected template. We find that, in the considered examples, the memory amplitude can still be effectively constrained compared to the case without correction. Soft signals may also contribute to the stochastic gravitational-wave background. We simulate an idealized background of soft displacement-memory signals, and evaluate its detectability at the considered detectors. Our results highlight soft signals from bursts with memory as potential targets for space-based gravitational-wave detectors.
@@ II Soft memory waveforms
The GWs emitted by distant sources are well-approximated by plane waves when they reach our detectors. Consider a flat spacetime perturbed by a weak plane GW in the transverse-traceless (TT) gauge:
@@
To see the intuitive observable effects of displacement and velocity memory, we introduce a free-falling timelike test particle at $\mathbf{x=0}$ in the spacetime (1) and the Fermi normal coordinates (FNC) $\bar{x}^{a}=(\bar{t},\bar{\mathbf{x}})$ centered at it, with $\frac{\partial x^{\mu}(\bar{x})}{\partial\bar{x}^{a}}\approx\delta^{\mu}_{a}$ and the particle’s proper time $\bar{t}\approx t$. In this FNC, the equation of motion for a nearby free-falling classical test particle with $\bar{x}^{j}(\bar{t}=0)=X^{j}$ and negligible initial velocity is
@@
The response of an interferometer to low-frequency GWs can be understood as arising from such relative displacements between test masses in the FNC, where the effects of GWs on light propagation are negligible. However, this approximation is not adequate for space-based detectors, and (as discussed in Sec. III) the signals must be analyzed in TT coordinates.
@@ II.1 Displacement memory
Displacement memory refers to a DC (direct current) shift of $h(t)$ between the asymptotic past and future, i.e., $\Delta h\equiv h(\infty)-h(-\infty)\neq 0$. In Eq. (6), this leads to a “permanent” displacement: $\Delta\bar{x}^{i}\approx\frac{1}{2}\Delta h_{ij}X^{j}$. In practice, $\infty$ refers to a moment that is so much far away from the observer’s time of interest. The soft waveform of displacement memory depends solely on the height $\Delta h$ of the memory jump. As illustrative examples, we consider three simple models with exactly computable frequency spectra. The first one is the arctan model:
@@
with $f_{*}>0$. Note that here for the Fourier transform we are treating $h(t)$ as a tempered distribution, and a global DC shift of the time-domain waveform is unobservable in the spectrum at $f>0$. The jump arises exactly from the zero-frequency mode, which can be seen by imposing an infrared cutoff at $f_{c}>0$. For the waveform (8), one has
$\int_{f_{c}}^{\infty}df\,e^{-2\pi ift}\tilde{h}_{\infty}(f)=\frac{i\Delta h}{2\pi}\Gamma\left(0,2\pi f_{c}(if_{*}t+1)/f_{*}\right),$ (9)
@@
In all three models above, $1/f_{*}$ measures the time scale of the jump. In the limit $f_{*}\to\infty$, the waveform tends to a step at $t=t_{*}$, with the height $\Delta h$. For a finite $f_{*}$, the models have the same zero-frequency limit at $f/f_{*}\to 0^{+}$:
@@
since the details of the jump are unresolvable in this limit. Note that the spectral density of the graviton number flux associated with this waveform diverges at zero frequency, but that of the energy flux remains finite. (10) is shared by all waveforms with the same displacement memory $\Delta h$. For $f\ll 1/t_{*}$, the phase factor $e^{2\pi ift_{*}}$ can be further approximated by unity. This however is generally a bad approximation for the GW signal, since the frequency spectrum of the waveform can only be measured at finite frequency, with the lowest resolvable frequency limited by the signal duration and instrumental noise. By the same reason, detection of the soft waveform (10) cannot directly prove the existence of the memory jump (which is associated with zero-frequency gravitons), but will provide a decisive evidence. Likewise, a detector capable of observing the spectrum in the form (10) at frequency $f>f_{\text{min}}$ has no information about it at lower frequencies. A more complete measurement of the waveform can thus only rely on the detectors having better sensitivities at low frequencies.
We note that for a waveform $h(t)$ with displacement memory satisfying $h(t<t_{\text{min}})\approx h(-\infty)$ and $h(t>t_{\text{max}})\approx h(\infty)$, $\tilde{h}_{\infty}(f)$ is drastically different from the Fourier transform of $h(t)$ in the finite interval $t_{\text{min}},t_{\text{max}}]$:
$\tilde{h}(f)$ $\equiv\int_{t_{\text{min}}}^{t_{\text{max}}}dt\,e^{2\pi ift}\,h(t)$
$=\tilde{h}_{\infty}-\frac{i}{2\pi f}\left[h(\infty)\,e^{2\pi ift_{\text{max}}}-h(-\infty)\,e^{2\pi ift_{\text{min}}}\right],$
which in particular is dominated by the second term at high frequencies. For a GW interferometer, however, the observable is not $h(t)$, but its temporal difference: h(tD)h(t)h(t-D)-h(t), with $D$ being constant within a sufficiently short time. So long as $h(t_{\text{min}}+\text{min}(0,-D))\approx h(-\infty)$ and $h(t_{\text{max}}+\text{max}(0,-D)\approx h(\infty)$, because
$\int_{t_{\text{min}}}^{t_{\text{max}}}dt\,e^{2\pi ift}[h(t-D)-h(t)]$ (11)
$\qquad\approx\int_{-\infty}^{\infty}dt\,e^{2\pi ift}[h(t-D)-h(t)]$
$\qquad=\left(e^{2\pi ifD}-1\right)\tilde{h}_{\infty}(f),$
the frequency spectrum of the signal well below the sampling frequency can be approximately derived with $\tilde{h}_{\infty}(f)$, but not $\tilde{h}(f)$. Numerically, $\tilde{h}_{\infty}(f)$ can be simply obtained from $\dot{h}(t)$ as $\tilde{\dot{h}}_{\infty}/(-2\pi if)$, with $\tilde{\dot{h}}_{\infty}\approx\tilde{\dot{h}}$ well-approximated by its discrete Fourier transform (DFT). Another way of computing $\tilde{h}_{\infty}(f)$ is to decompose $h(t)$ as a simple waveform with memory such as the arctan model, and a remaining part that has no memory; the latter is then well-approximated by its DFT.
@@
If the time interval $t_{\text{min}},t_{\text{max}}]$ is too short to include the entire jump, the observed signal will be quite different from the soft waveform. As an extreme example, consider a linearly growing waveform $\dot{h}(t)=\text{const}$, for which $h(t-D)-h(t)\propto D$ is approximately constant, and will actually be removed (as a DC component) in the data processing. Moreover, in this paper, we focus on the measurements using space-based interferometers. As explained in Sec. III, the relevant time-domain signals (e.g., Eq. (38)) are built up by observables such as $h(t-D_{1})-h(t)]-[h(t-D_{1}-D_{2})-h(t-D_{2})]$, with $D_{1,2}$ being constant within a sufficiently short time. A linearly growing waveform thus has a vanishing signal in the static-detector approximation. In fact, we find that the signal remains extremely weak compared with that of a soft memory waveform with $\Delta h=\dot{h}\,T$ when the signal duration $T$ is not short, so that the static-detector approximation no longer holds.

### 13 Apr 2026 v2

@@ Abstract
Gravitational-wave displacement memory is a remarkable and ubiquitous phenomenon predicted by general relativity, which has not yet been detected. Unlike the oscillatory components of gravitational waveforms, displacement memory is associated with soft gravitons, making it the only observable signal of its parent event at sufficiently low frequencies. Similarly, soft waveforms may arise from velocity and integrated-displacement memory. The simple and universal spectral shapes of soft waveforms also provide effective templates for matched filtering and parameter estimation. In this paper, we investigate the detection prospects for such soft memory signals with future space-based laser interferometers. As realistic examples, we examine the infrared spectral features of gravitational waves from moderately relativistic compact binary scattering and nearly equal-mass quasi-circular, non-precessing black hole mergers. In both cases, the low-frequency spectrum can be described by a corrected soft waveform of displacement memory. The results of simulated Bayesian parameter estimation demonstrate that independent measurement of a soft displacement-memory signal with a single LISA-like detector is achievable at signal-to-noise ratios $\gtrsim 10$. The measurement precision can be significantly improved by joint observations with a LISA-Taiji network. A single BBO detector could be capable of separately measuring the null memory from stellar-mass compact binary mergers. We also evaluate the detectability of an idealized stochastic background of soft displacement-memory signals. Our results indicate that gravitational-wave bursts with memory can be promising targets for space-based interferometers.
@@ I Introduction
Besides the oscillatory components, gravitational radiation in general relativity generally contains a non-oscillatory component that results in a displacement memory [12, 21, 92, 34]. It may have an “ordinary” or “null” origin [10], and is associated with the supertranslation charge balance laws of asymptotically flat spacetimes [98, 36, 85, 72] as well as with Weinberg’s soft graviton theorem [85]. Displacement memory can therefore serve as an important test of general relativity. In addition to their theoretical interest, gravitational wave (GW) bursts with displacement memory are expected to be produced by a variety of astrophysical sources, including compact binaries [33, 32, 35, 91, 101, 29, 94, 68, 107, 60, 52, 103, 27, 49, 69, 42, 53, 31, 25, 7], gamma-ray bursts [81, 82], and core-collapse supernovae [73, 77]. Searches for displacement memory in the GWs from compact binary mergers have been considered for ground-based detectors [67, 71, 101, 30, 58, 57, 94, 47, 20, 45], space-based detectors [61, 44, 43, 59, 88, 45, 22, 108], and pulsar timing arrays [95, 96, 61, 3, 86, 28, 87, 2]. Potentially observable imprints of displacement memory may also appear in the cosmic microwave background [109].
@@
Although displacement memory ubiquitously occurs in gravitational radiation, it has so far remained undetected. For example, the null memory of a compact binary merger produces only a small modification of the waveform and is not easily disentangled from the oscillatory “parent” signal in the time domain. Resolving such modifications through sensitive detectors and refined waveform modeling would enable the detection of memory and the improvement of parameter estimation [47, 43, 59, 88]. Accurately modeling waveforms that include memory effects becomes increasingly challenging for more complex sources. For a compact binary scattering event, even with a sufficiently accurate waveform model, parameter estimation is still challenging due to degeneracies [37]. Waveform modeling is further complicated by possible environmental influences [18, 84, 4]. However, because of its soft nature, displacement memory produces a signal that persists at low frequencies, with a spectrum of universal shape. Although much information about the sources is lost in such “orphan” signals [95, 71, 44], they remain valuable targets for probing the memory components—either by searching for them or by measuring them independently—thanks to the ubiquity of the sources and the simplicity of the templates. It is also possible that some yet-unknown processes could produce observable GW bursts with displacement memory, making soft memory signals a potential probe of these processes.
@@
In addition to the soft waveform associated with displacement memory (which corresponds to a step-like strain in the time domain), other types of soft waveforms may also be observed, including those associated with velocity memory [48, 11] (corresponding to a step-like first derivative of strain in the time domain) and integrated-displacement memory (discussed in Sec. II.3). Although these waveforms are not predicted by the soft theorem, they may arise over sufficiently long time scales and, in such a limit, serve as useful descriptions.
@@
In this paper, we study the prospects for detecting and measuring the soft memory signals with future space-based interferometers, including Laser Interferometer Space Antenna (LISA) [5, 6], Taiji [55], TianQin [55], and Big Bang Observer (BBO) Stage I [23]. We investigate the detector responses and the parameter estimation precision for both a single LISA-like detector and the proposed LISA-Taiji network [80]. In optimal configurations, the signal-to-noise ratio in LISA, Taiji and TianQin can reach 10 for a displacement memory of amplitude $\sim 10^{-20}$, or a velocity memory of amplitude $\sim 10^{-22}\,\text{Hz}$. To examine the precision of the measurement, we perform numerical simulations for Bayesian parameter estimation. We find that with a single detector, the memory parameters may be effectively probed for signal-to-noise ratios $\gtrsim 10$, although the measurement precision is limited by the signal’s degeneracy at low frequencies. This limitation can be mitigated by using a detector network or by incorporating prior information on the source direction. We demonstrate that the joint observations by LISA and Taiji can achieve significantly more precise measurements. Soft memory signals in BBO (occurring at $f\lesssim 1\,\text{Hz}$) appear to suffer from less degeneracy and can achieve much higher sensitivity, with the signal-to-noise ratio reaching 10 for a displacement memory of amplitude $\sim 10^{-24}$, or a velocity memory of amplitude $\sim 10^{-23}\,\text{Hz}$. This makes it possible to independently measure the null displacement memory from stellar-mass compact binary mergers.
@@
The soft waveform is an idealized zero-frequency limit. We examine the correction to the soft waveform in two realistic examples: hyperbolic encounter of compact binary and nearly equal-mass quasi-circular binary black hole (BBH) merger, and assess its impact on displacement memory measurements using the uncorrected template. We find that, in the considered examples, the memory amplitude can still be effectively constrained compared to the case without correction. Soft signals may also contribute to the stochastic gravitational-wave background. We simulate an idealized background of soft displacement-memory signals, and evaluate its detectability at the considered detectors. Our results highlight soft signals from bursts with memory as potential targets for space-based gravitational-wave detectors.
@@ II Soft memory waveforms
The GWs emitted by distant sources are well-approximated by plane waves when they reach our detectors. Consider a flat spacetime perturbed by a weak plane GW [^1] The memory effects under discussion should not be confused with those in exact pp-wave spacetimes [1]. in the transverse-traceless (TT) gauge:
@@
To illustrate the observable effects of displacement and velocity memory, we introduce a free-falling timelike test particle at $\mathbf{x=0}$ in the spacetime (1) and the Fermi normal coordinates (FNC) $\bar{x}^{a}=(\bar{t},\bar{\mathbf{x}})$ centered at it, with $\frac{\partial x^{\mu}(\bar{x})}{\partial\bar{x}^{a}}\approx\delta^{\mu}_{a}$ and the particle’s proper time $\bar{t}\approx t$. In this FNC, the equation of motion for a nearby free-falling classical test particle with $\bar{x}^{j}(\bar{t}=0)=X^{j}$ and negligible initial velocity is
@@
The response of an interferometer to low-frequency GWs can be understood as arising from such relative displacements between test masses in the FNC, where the effects of GWs on light propagation are negligible. However, this approximation is not adequate for space-based detectors, and (as discussed in Sec. III) it is convenient to analyze the signals in the TT coordinates.
@@ II.1 Displacement memory
Displacement memory refers to a DC (direct current) shift of $h(t)$ between the asymptotic past and future, i.e., $\Delta h\equiv h(\infty)-h(-\infty)\neq 0$. In Eq. (6), this leads to a “permanent” displacement: $\Delta\bar{x}^{i}\approx\frac{1}{2}\Delta h_{ij}X^{j}$. In practice, $\infty$ refers to a moment that is so much far away from the observer’s time of interest. The soft waveform of displacement memory depends solely on the amplitude $\Delta h$ of the memory jump. As illustrative examples, we consider three simple models with exactly computable frequency spectra. The first one is the arctan model:
@@
with $f_{*}>0$. Note that here for the Fourier transform we are treating $h(t)$ as a tempered distribution, and a global DC shift of the time-domain waveform is unobservable in the spectrum at $f>0$. The asymptotic transition is due to the zero-frequency mode, which can be seen by imposing an infrared cutoff at $f_{c}>0$. For the waveform (8), one has
$\int_{f_{c}}^{\infty}df\,e^{-2\pi ift}\tilde{h}_{\infty}(f)=\frac{i\Delta h}{2\pi}\Gamma\left(0,2\pi f_{c}(if_{*}t+1)/f_{*}\right),$
@@
In all three models above, $1/f_{*}$ measures the time scale of the jump. In the limit $f_{*}\to\infty$, the waveform approaches a step function at $t=t_{*}$ with the step size $\Delta h$, corresponding to $\tilde{h}_{\infty}(f)=[i\Delta h/(2\pi f)+\delta(f)/2]e^{2\pi ift_{*}}$. For a finite $f_{*}$, the models have the same zero-frequency limit at $f/f_{*}\to 0^{+}$:
@@
since the details of the jump are unresolvable in this limit. Note that the spectral density of the graviton number flux, $dN/(dAdf)=[(\pi/2)f^{2}|\tilde{h}_{\infty}|^{2}]/(2\pi f)$, associated with this waveform diverges at zero frequency, whereas the spectral density of the energy flux remains finite. (9) is shared by all waveforms with the same displacement memory $\Delta h$. For $f\ll 1/t_{*}$, the phase factor $e^{2\pi ift_{*}}$ can be further approximated by unity. This however is generally a bad approximation for the GW signal, since the frequency spectrum of the waveform can only be measured at finite frequency, with the lowest resolvable frequency limited by the signal duration $t_{\text{min}},t_{\text{max}}]$ and instrumental noise. For the same reason, detection of the soft waveform (9) cannot demonstrate the persistence of the memory jump, although it would provide decisive evidence. Likewise, even if a detector observes the spectrum in the form (9) at $f>f_{\text{min}}$, this does not determine the spectrum at lower frequencies. A more complete measurement of the waveform therefore requires detectors with better sensitivity in the low-frequency band.
We note that for a waveform $h(t)$ with displacement memory satisfying $h(t<t_{\text{min}})=h(-\infty)$ and $h(t>t_{\text{max}})=h(\infty)$, $\tilde{h}_{\infty}(f)$ is drastically different from the Fourier transform of $h(t)$ in the finite interval $t_{\text{min}},t_{\text{max}}]$: $\tilde{h}(f)\equiv\int_{t_{\text{min}}}^{t_{\text{max}}}dt\,e^{2\pi ift}\,h(t)=\tilde{h}_{\infty}-\frac{i}{2\pi f}\left[h(\infty)\,e^{2\pi ift_{\text{max}}}-h(-\infty)\,e^{2\pi ift_{\text{min}}}\right]$, which in particular is dominated by the second term at high frequencies. For a GW interferometer, however, the observable is not $h(t)$, but its temporal difference: h(tD)h(t)h(t-D)-h(t), with $0<D\ll|t_{\text{max}}-t_{\text{min}}|$ approximately constant over the signal duration. If $h(t_{\text{max}}-D)=h(\infty)$, since $\int_{t_{\text{min}}}^{t_{\text{max}}}dt\,e^{2\pi ift}[h(t-D)-h(t)]=\int_{-\infty}^{\infty}dt\,e^{2\pi ift}[h(t-D)-h(t)]=\left(e^{2\pi ifD}-1\right)\tilde{h}_{\infty}(f)$, the frequency spectrum of the signal well below the sampling frequency can be approximately derived with $\tilde{h}_{\infty}(f)$, but not $\tilde{h}(f)$. Numerically, $\tilde{h}_{\infty}(f)$ can be obtained from $\dot{h}(t)$ as $\tilde{\dot{h}}/(-2\pi if)$, with $\tilde{\dot{h}}$ well-approximated by its discrete Fourier transform (DFT). Even if the memory jump does not occur strictly within $t_{\text{min}},t_{\text{max}}]$, the approximation typically remains valid for $f>1/|t_{\text{max}}-t_{\text{min}}|$, provided that $h(t)$ varies slowly outside this interval, in which case $\tilde{\dot{h}}/(-2\pi if)$ is close to $\tilde{h}_{\infty}(f)$. In subsequent sections, we compute the numerical Fourier spectrum of a time-domain waveform $h(t)$ over a finite interval $t_{\text{min}},t_{\text{max}}]$ as $\int_{t_{\text{min}}}^{t_{\text{max}}}dt\,e^{2\pi ift}\,\dot{h}(t)\,/(-2\pi if)$.
@@
If the interval $t_{\text{min}},t_{\text{max}}]$ is too short to capture most of the jump, the signal spectrum at $f\gtrsim 1/|t_{\text{max}}-t_{\text{min}}|$ can differ significantly from that of the soft waveform. As an extreme example, consider a linearly growing waveform $\dot{h}(t)=\text{const}$, for which $h(t-D)-h(t)\propto D$ is approximately constant, and will actually be removed (as a DC component) in the data processing. Moreover, in this paper, we focus on the measurements using space-based interferometers. As explained in Sec. III, the relevant time-domain signals (e.g., Eq. (31)) are built up by observables such as $h(t-D_{1})-h(t)]-[h(t-D_{1}-D_{2})-h(t-D_{2})]$, with $D_{1,2}$ being constant within a sufficiently short time. A linearly growing waveform thus has a vanishing signal in the static-detector approximation. In fact, we find that the signal remains extremely weak compared with that of a soft waveform with $\Delta h=\dot{h}\,T$ when the signal duration $T$ is not short, so that the static-detector approximation no longer holds.